# Clustering by Mining Density Distributions and Splitting Manifold Structure

### v1

@@ 3. $c$:
Fig. 5 evaluates the performance variation of the algorithm under different hyperparameter settings, we tested it on multiple datasets. shows the ARI variations of our algorithm, SC, and DPC-DBFN algorithms on datasets optical_test, Pendigits, Kdd9, and mfea-fac. The results indicate that our algorithm experiences larger fluctuations with smaller $k$ values on the Pendigits and Kdd9 datasets, but the ARI values stabilize as $k$ increases. Furthermore, our algorithm is relatively stable on the optical_test and mfeafac datasets. This suggests that our algorithm exhibits a certain degree of robustness to changes in $k$ values and maintains high performance across most datasets.
@@ Conclusion
The paper proposes a new approach to represent data in a coarse granularity that can accelerate and improve the performance of Spectral Clustering. Specifically, the approach can discover micro-clusters based on local density distributions of data, and introduces the concept of manifold curvature of micro-clusters to help split them into more convex ones. In this way, the approach provides better representation of data and simplifies the characterization of the similarities, resulting in better performance in subsequent Spectral Clustering. Evaluations on 4 synthetic datasets and 13 real-world datasets, against relevant state-of-the-art algorithms, show that the proposed algorithm performs best on most datasets. The ablation experiments also demonstrated the effectiveness of the proposed components.
Despite the impressive performance of the proposed algorithm in the experiments, there is still room for improvement and optimization. Future work will attempt to introduce more efficient computational techniques, such as parallel and distributed computing, to further enhance the computational speed of the algorithm, especially when handling ultra-large-scale datasets. Additionally, one can consider developing adaptive hyperparameter optimization methods for this approach, and can also consider to combine it with other clustering methods.

### v2

@@ 3. $c$:
Fig. 6 demonstrates the ARI variations of MDMSC under different hyperparameter settings on multiple datasets, where $k$ was set in $[3,50]$, $\lambda$ in $[1.0,3.0]$ with a step size of 0.2, and $\beta$ in $[8,16]$. From the results in Fig. 6, it is evident that while the algorithm exhibits some fluctuations with respect to $k$ (larger fluctuations with smaller $k$ and stablized as $k$ increases), its performance remains relatively stable when varying $\lambda$ and $\beta$. This suggests that MDMSC has a certain degree of robustness against changes in these parameters.
@@ Conclusion
The paper proposes a new approach to represent data in a coarse granularity that can accelerate and improve the performance of spectral clustering. Specifically, the approach can discover micro-clusters based on local density distributions of data, and introduces the concept of manifold curvature of micro-clusters to help split them into more convex ones. In this way, the approach provides better representation of data and simplifies the characterization of the similarities, resulting in better performance in subsequent spectral clustering. Evaluations on 4 synthetic datasets and 13 real-world datasets, against relevant state-of-the-art algorithms, show that the proposed algorithm performs best on most datasets. The ablation experiments also demonstrate the effectiveness of the proposed components.
There is still room for improvement and optimization. Future work will attempt to introduce more efficient computational techniques, such as parallel and distributed computing, to further enhance the computational speed of the algorithm, especially when handling ultra-large-scale datasets. Additionally, one can consider developing adaptive hyperparameter optimization methods for this approach, and can also consider to combine it with other clustering methods.
Acknowledgments
We would like to thank the anonymous reviewers for their invaluable help to improve the paper. This work was supported by the National Natural Science Foundation of China [Nos. 61806170 and 62276218], the Fundamental Research Funds for the Central Universities [Nos. 2682022ZTPY082 and 2682023ZTPY027], and the Open Foundation of Key Laboratory of Cyberspace Security, Ministry of Education of China and Henan Key Laboratory of Cyberspace Situation Awareness [No.KLCS20240105].